# A Spatial-statistical model to analyse historical rutting data

### v2

@@ Abstract
Pavement rutting poses a significant challenge in flexible pavements, necessitating costly asphalt resurfacing. To address this issue comprehensively, we propose an advanced Bayesian hierarchical framework of latent Gaussian models with spatial components. Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates. Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), pavement type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting. This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year). We apply our models to a data set spanning eleven years (2010-2020). Our findings emphasize the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components. Pavement type, in conjunction with traffic intensity, is also found to be the primary driver of rutting. Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations. This leads to a halving of the expected pavement lifespan in these areas. Our study offers valuable insights, presenting maps indicating expected rutting, and identifying locations with accelerated rutting rates, resulting in a reduction in pavement life expectancy of at least 10 years.
@@ 1 Introduction
Several empirical models have been proposed in the past to predict the performance of pavements (Hu et al. 2022, Yao et al. 2023, Yu et al. 2023) and these models typically include rutting. Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., pavement or asphalt type, weather, width of the road, layer thickness). Empirical models are known to suffer from limited applicability (Gupta et al. 2014), so when the conditions (for example road construction, weather conditions or damage mechanisms) during the data collection differs from the conditions where the model is applied, one cannot expect to get reasonable results. Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present. Calculating the expected rutting for sections based on an empirical model is therefore likely to give large uncertainty.
@@
Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced spatial rutting effects. To achieve this, we employ a hierarchical Bayesian framework comprising latent Gaussian models (LGMs) that encompass both random and spatial components (Jourdain and Klein-Paste 2024). Latent Gaussian models have been used in various statistical analyses, as demonstrated by their applications in a multitude of fields (e.g., Dey et al. 2000, Chu et al. 2005, Banerjee et al. 2008, Rue et al. 2009b). These models offer a high degree of flexibility and explicitly account for model dependence. They can approximate entire processes defined over continuous domains and provide probabilistic predictions (Lindgren and Rue 2015). In our specific case, the rutting phenomenon on road surfaces persists regardless of whether it is directly observed, with the latent fields serving as the authentic representation of this rutting. A crucial premise underlying spatial models is that observations situated in close proximity to one another in the spatial domain exhibit greater similarity than those positioned further apart. For a broad class of LGMs, the Integrated Nested Laplace Approximation approach (INLA, Rue et al. 2009a) proves to be an apt tool for performing Bayesian inference.
The proposed models are based on historical rutting observations and a selection of explanatory variables known to influence pavement rutting that is available for the Norwegian stretch of the road. These variables are pavement type, Average Annual Daily Traffic (AADT), rut depth from the previous year and lane width. The inclusion of these variables is grounded in existing knowledge of factors that significantly influence rutting in road surfaces (Chan et al. 2010, Lang 2012, Thodesen et al. 2012, Ebrahimi et al. 2019, Yan et al. 2020, Zhao et al. 2020, Pan et al. 2021, Li et al. 2023). A critical element in our model is the role of Average Annual Daily Traffic (AADT) on the pavement. Rutting, as we argue, is contingent on the presence of AADT or the load imposed by traffic. Without traffic loading, certain forms of pavement deterioration, such as cracks and roughness, may occur, but rutting is improbable. This hypothesis aligns with prior research on the prediction of pavement variability and deterioration, as illustrated in the work of Hao et al. (2009) and Vedvik (2021). Rutting is a complex phenomenon, and the depth of rutting observed in the previous year is a key factor in predicting future rutting (Archilla and Madanat 2000). Deeper rut depths from the previous year are expected to correlate with increased rutting in the current year, and this relationship is incorporated into the model. Additionally, lane width is considered as a contributing factor to rutting rates. It is logical to assume that narrower roads are more prone to higher rutting rates due to the limited space for motorists to maneuver. However, it’s important to note that the effect of lane width on rutting is contingent on the intensity of traffic. Pavement type also plays a crucial role, particularly in colder climates where it can significantly influence rutting resistance. During winter, rising temperatures lead to shorter periods of frozen ground and snow cover on road surfaces, altering the frequency of freeze-thaw cycles. Consequently, this poses challenges for lower-grade binder types like asphalt gravel concrete (Agc), which are more susceptible to rutting under such conditions
As proof of concept, our proposed models are applied to eleven years (2010-2020) of pavement rutting data for one road stretch; the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border. We identify areas with unexplained or accelerated rutting and determine which explanatory variables are most important to rutting development. A detailed description of the data used and data cleaning process is given in Section 2. Model notation, formulation and evaluation criteria are given in Section 3. Results and discussion are given in Section 4. Section 5 explains how road authorities can use the model, based on the results.
@@ 2 Materials
2.1 Data source
The data material used in this research is obtained from the Norwegian Road Database (NVDB) and the Rosita database. The NVDB is primarily developed by the Norwegian Public Roads Administration (NPRA), the road authority for national roads, including managing, researching, planning, building, and operating. In 2022, NVDB consisted of 93 500 kilometres of public roads, including national roads (state-owned, 10,500 kilometres), county roads (44 000 kilometres) and municipal roads (39 000 kilometres) (Stephansen 2022). The Rosita database, which stores extensive and calculated data from measuring systems (see Section 2.3), was made available for use in this study.
@@ 2.3 Data
The available explanatory variables for expected rut depth includes, traffic load in terms of annual average daily traffic (AADT), pavement or asphalt type (three categories), rut depth from the previous year and the lane width. A summary of the data used can be found in Table 1 and Figure 1 and Supplementary Material LABEL:supplementary:datasummary.
@@
Terms and Variables Description/Value Rut depth ($d_{i,t}$, millimetre) Measurement for year $t$ for all $20$-metre stretch ($i$); mean $=12.11$ millimetres and sd $=6.78$ millimetres Lane width (${w_{i}}$, metre) Half (or quarter) of width of 2-lane (or $2-2$ lane) used; mean $=6.02$ millimetres and sd $=0.69$ millimetres Pavement type Asphalt concrete (Ac), asphalt gravel concrete (Agc) and stone mastic asphalt (Sma) AADT (vehicles/year) $568-18500$ vehicles, with half (or quarter) of AADT for 2-lane (or $2-2$ lane) road used $r_{i,t}$ (millimetre/year) Change in rut depth between year $t$ and $t-1$ (for $t=1,...,T$).
@@ 2.3.1 Rut depth and rutting
The rut depth is measured annually using a laser scanner device ViaPPS (Pavement Profile Scanner) — a measurement system developed by ViaTech in cooperation with NPRA (Peraka and Biligiri 2020). The annual rut depth measurement is used to calculate rutting (mm/year), the indicator variable used for assessing pavement performance. We denote the rut depth variable at road segment $i=1,...,n$ in year $t$ as $d_{i,t}$, and $r_{i,t}$ as rutting (mm/year). Rutting for year $t$ at segment $i$ is defined as
@@
In the case of repavement, rut depth is smaller in the proceeding year resulting in large negative values of $r_{i,t}$. Therefore, following Vedvik (2021), large negative rutting values are removed;
$r_{i,t}=\begin{cases}r_{i,t},&\text{if }r_{i,t}\geq-d_{i,t}/2\\ \text{NaN }&\text{otherwise, }\end{cases}$ (2)
i.e., $r_{i,t}$ is kept if it is at least half the rut depth, or assigned NaN (not a number) otherwise. Some missing rutting data may occur due to uncertainty arising from the actual rut depth measurements or the data collection process (Figure 1, graph (c) shows large variability in rutting).
@@ 2.3.2 Annual Average Daily Traffic (AADT)
The annual average number of daily traffic (AADT) f is available for each road segment for each year. AADT ranges from 568-18500 between the years $2010-2020$, with more traffic close to Stjørdal and less towards Sweden. NPRA states that high-traffic roads have an AADT of more than 5000 vehicles and recommends maintenance at rut depth values exceeding $20$ millimetres, while low-traffic roads with AADT lower than 5000 require pavement treatment when rut depth exceeds $25$ millimetres. Up to $2.5$ kilometres along the road, a quarter of the AADT is used in the analysis or half of the AADT otherwise. In all analyses, AADT in ten thousands are used.
2.3.3 Pavement type
In Norway, asphalt concrete is used on all types of roads and in all traffic classes. The asphalt mixtures used on the European route E14 are; asphalt concrete mixtures (Ac) and asphalt gravel concrete (Agc), and a stone mastic asphalt mixture (Sma). These include stone mastic asphalt Sma is the stronger and densely graded asphalt mass with unmodified binders and is mainly used on high-traffic roads; and asphalt concrete (Ac) and asphalt gravel concrete (Agc), which are mainly used on low-traffic roads of up to 5000 vehicles, courtyards, sidewalks and bicycle paths. These are softer, dynamic pavement mixtures, with less stringent requirements for the nature and grading of the stone material compared with that of stone mastic asphalt. Asphalt concrete, is however, more stringent that Asphalt gravel concrete; it is a dense-graded asphalt used as a protective layer over the road surface, or laid over a layer of asphalt gravel concrete (d’Angelo et al. 2008, vegvesen Vegdirektoratet 2009, Taddesse 2013, Zhang et al. 2022).
2.3.4 Lane width
The road width is quite variable and depends on the number of lanes. Since one lane is considered for the purposes of analysis, a quarter of the measured width up to 2.5 kilometres from Stjørdal or half of the measured width for the rest of the road stretch is used.
(a) AADT/Pavement type, and lane width
@@ 2.3.4 Lane width
Figure 1: AADT together with pavement type, and lane width (a), and rutting (for years 2019-2020, mm/year — missing in red) (b).
@@ 2.4 Exploratory analysis of rutting
On average, over all stretches and years, rutting is $1.62$ mm/year; see Figure  2 (a). The density plots,  2 graph (b), show rutting for all segments for each year. Some years having higher rutting levels and are more widely distributed than others. The lines in graphs (c) and (d) give the correlation of rutting values with respect to distance (mean across all years, and two randomly selected years; 2011 and 2012, respectively) (Bachmaier and Backes 2011). It is computed from the observed rutting and shows how the rutting values are correlated with respect to distance between the corresponding segments. The distance (metre) is shown on the x-axis, and the calculated correlation between rutting values is on the y-axis. The graph shows some spatial dependencies of pavement rutting along the road. These exist for short distances of up to ($\sim$) $300$ metres (3 kilometres).
Refer to caption (a) Observed rutting Refer to caption (b) Density Refer to caption (c) Autocorrelation across years Refer to caption (d) Autocorrelation for years 2011-2012
Figure 2: Graph (a): Observed rutting for each year (2011-2020), observed mean rutting $1.62$ mm/year (     ), and the observed mean rutting across segments ( ). Graph (b): Rutting density plots, graph. Graphs (c) and (d) show how rutting values are correlated with distance. Spatial dependencies in rutting exist up to 3 kilometres.
@@ 3 Latent Gaussian models for Rutting
A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters. In this paper, the true rutting at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects. We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
3.1 Likelihood: observed rutting given true rutting
We denote the observed rutting for segment $i$ in year $t$ as $r_{i,t}$. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as

### v3

# interacttfvsample_Supplementary

@@ Abstract
Pavement rutting poses a significant challenge in flexible pavements, necessitating costly asphalt resurfacing. To address this issue comprehensively, we propose an advanced Bayesian hierarchical framework of latent Gaussian models with spatial components. Our model provides a thorough diagnostic analysis, pinpointing areas exhibiting unexpectedly high rutting rates. Incorporating spatial and random components, and important explanatory variables like annual average daily traffic (traffic intensity), asphalt type, rut depth and lane width, our proposed models account for and estimate the influence of these variables on rutting. This approach not only quantifies uncertainties and discerns locations at the highest risk of requiring maintenance, but also uncover spatial dependencies in rutting (millimetre/year). We apply our models to a data set spanning eleven years (2010-2020). Our findings emphasise the systematic unexplained spatial rutting effect, where some of the rutting variability is accounted for by spatial components, asphalt type, in conjunction with traffic intensity, is also found to be the primary driver of rutting. Furthermore, the spatial dependencies uncovered reveal road sections experiencing more than 1 millimeter of rutting beyond annual expectations. This leads to a halving of the expected pavement lifespan in these areas. Our study offers valuable insights, presenting maps indicating expected rutting, and identifying locations with accelerated rutting rates, resulting in a reduction in pavement life expectancy of at least 10 years.
@@ 1 Introduction
Several empirical models have been proposed in the past to predict the performance of pavements (Hu et al. 2022, Yao et al. 2023, Yu et al. 2023) and these models typically include rutting. Empirical models usually relate the accumulation of distresses (e.g., rut, roughness, cracks) in the pavement layers to the load repetitions (annual average daily traffic) and other affecting factors (e.g., asphalt type, weather, width of the road, layer thickness). Empirical models are known to suffer from limited applicability (Gupta et al. 2014), so when the conditions (for example road construction, weather conditions or damage mechanisms) during the data collection differs from the conditions where the model is applied, one cannot expect to get reasonable results. Empirical models developed for pavement wear in cold climates exists (Jacobson and Wågberg 2005), but the challenge of limited predictability remains present. Calculating the expected rutting for sections based on an empirical model is therefore likely to give large uncertainty.
@@
1.1 Objectives
Our primary goal is to develop spatial-statistical models that can pinpoint pavement locations exhibiting unexpected and pronounced rutting. To achieve this, we use eleven years (2010-2020) of pavement rutting data from the European highway route E14, stretching $67.1$ kilometres from Stjørdal, Norway to Storlien on the Swedish border. This data, collected annually, provides a robust historical record that is essential for identifying trends and anomalies in rutting patterns. By utilising this extensive dataset, we aim to enhance our understanding of rutting dynamics over time and across different sections of the road.
To analyze the rutting phenomenon, we develop a hierarchical Bayesian framework utilizing latent Gaussian models (LGMs) (Jourdain et al. 2024). These models are particularly suited for this type of analysis due to their flexibility and ability to explicitly account for spatial and random components (e.g., Dey et al. 2000, Chu et al. 2005, Banerjee et al. 2008, Rue et al. 2009b). The LGMs enable us to capture the underlying processes governing rutting and provide probabilistic predictions, which are crucial for understanding and managing pavement performance. Advanced statistical analysis is implemented through the Integrated Nested Laplace Approximation approach (INLA) for performing Bayesian inference. This approach is well-suited for LGMs and will allow us to efficiently handle the complex computations involved in our models. By using INLA, we aim to achieve high computational efficiency and accuracy in our analysis.
Incorporating a selection of explanatory variables known to influence pavement rutting is another key objective. These variables include asphalt type, Average Annual Daily Traffic (AADT), rut depth from the previous year, and lane width. The inclusion of these factors is based on their established impact on rutting (Chan et al. 2010, Lang 2012, Thodesen et al. 2012, Ebrahimi et al. 2019, Yan et al. 2020, Zhao et al. 2020, Pan et al. 2021, Li et al. 2023). By integrating these variables into our model, we can better account for the diverse factors that contribute to rutting, thereby improving the accuracy and reliability of our estimations.
Identifying areas with unexplained or accelerated rutting is a crucial part of our objectives. By analyzing the data with our models, we pinpoint sections of the road that exhibit unexpected rutting patterns.
Therefore, we provide practical insights for road authorities to use the model results for better pavement maintenance planning. By shifting from manual, suspicion-based diagnostics to a more efficient, automated approach, we hope to enable pavement managers to concentrate their diagnostic resources on identifying the correct countermeasures rather than merely locating problem areas. This will ultimately lead to more effective maintenance strategies and improved road safety.
@@ 2.1 Area of study
Most of the road has two lanes, but within approximately 2.5 km from Stjørdal, it expands to four lanes. Two-lane stretches are around 6 meters wide, while four-lane sections are up to 13 meters wide. The expected lifetime of the road is 20-30 years after construction.
The climate is subarctic or cold-temperate, with temperatures varying by season: winter (-2°C to 2°C), spring (5°C to 12°C), summer (15°C to 22°C), and autumn (10°C to 15°C) (Sygna and O’Brien 2001, Institutt 2010). On average, air surface temperatures are at most 18 °C (see Figure A.5 in Appendix A.5) .
For analysis, we focus on one lane (lane 1) in the direction from Stjørdal to Storlien.
@@ 2.2 Data
The data for this research is sourced from the Norwegian Road Database (NVDB) and the Rosita database. NVDB, primarily developed by the Norwegian Public Roads Administration (NPRA), includes 93,500 kilometers of public roads, comprising national, county, and municipal roads (Stephansen 2022). The Rosita database provides extensive and calculated data from measuring systems (see Section 2.2.1).
We use eleven years (2010-2020) of pavement rutting data from the European highway route E14, covering a 67.1-kilometer stretch from Stjørdal, Norway, to Storlien on the Swedish border. Collected annually, this data set offers a comprehensive historical record crucial for detecting trends and anomalies in rutting patterns.
Explanatory variables for expected rutting include annual average daily traffic (AADT), asphalt type, rut depth from the previous year, and lane width. Table 1 summarizes the data used, and Figure 1 and Supplementary Material LABEL:supplementary:datasummary provide additional details.
@@
Terms and Variables Description/Value
Rut depth ($d_{i,t}$, mm) Measurement for year $t$ for all $20$-metre road segments ($i$); mean $=12.11$, sd $=6.78$
Lane width (${w_{i}}$, m) Half (or quarter) of width of 2-lane (or $2-2$ lane) used; mean $=2.99$, sd $=0.18$
Asphalt type Asphalt concrete (Ac), asphalt gravel concrete (Agc), stone mastic asphalt (Sma)
AADT (vehicles/day) Range: 568-18,500, adjusted for the number of lanes
$r_{i,t}$ (mm/year) Rutting, calculated as the change in rut depth ($d_{i,t}$) between years $t$ and $t-1$
@@ 2.2.1 Rut Depth and Rutting
Rut depth is measured annually using the ViaPPS (Pavement Profile Scanner - laser scanner), developed by ViaTech in collaboration with the NPRA (Peraka and Biligiri 2020). Measurement positions are determined with high-accuracy GPS (CPOS/DPOS — satellite-based position service with centimetre/decimetre accuracy), ensuring precise location data. The rut depth is estimated by comparing the height of the ridges formed between the wheel paths, which involves identifying the two lowest points in the wheel path and measuring the distance to the top of the ridge. Additionally, the laser covers an approximate 4-meter footprint, ensuring that the vehicle remains within the wheel tracks and avoids crossing the centerline or shoulder. Thus, minor deviations (10-50 cm) do not affect the calculations. Annual rut depths along the road range from 0 to 41 mm, measured at 20-metre intervals. There are 3355 road segments measured for annual rut depth.
Rutting, denoted as $r_{i,t}$ (mm/year), is calculated as the difference in rut depth between years, as shown in Equation 1. We denote the rut depth variable at road segment $i=1,...,n$ in year $t$ as $d_{i,t}$, and $r_{i,t}$ as rutting (mm/year). Rutting for year $t=1,...,T$ at segment $i$ is defined as
@@
When repavement occurs, rut depth decreases, sometimes resulting in large negative values of the rutting ($r_{i,t}$). Following Vedvik (2021), large negative rutting values are considered missing and set to NaN (not a number):
$r_{i,t}=\begin{cases}r_{i,t}&\text{if }r_{i,t}\geq-\frac{d_{i,t}}{2}\\ \text{NaN}&\text{otherwise,}\end{cases}$
and the small negative values kept in the dataset are considered as measurement uncertainties and are accounted for in the random variability. Some missing rutting data may occur due to measurement uncertainty or data collection process. For available observations, the overall mean rutting and segment mean rutting (over all years) are calculated as $\bar{r}={1}/{n}\sum_{\forall t}\sum_{\forall i}r_{i,t}$ and $\bar{r}_{i}={1}/{T}\sum_{t=1}^{T}r_{i,t}$, respectively.
2.2.2 Lane Width
In the database, road width and the number of lanes are available. The first 2.5 km from Stjørdal has four (4) lanes, the remaining stretch has two lanes. Lane width is calculated by dividing the road by the number of lanes.
@@ 2.2.3 Annual Average Daily Traffic (AADT)
The Annual Average Daily Traffic (AADT) for each road segment between 2010 and 2020 varied from 568 to 18,500 vehicles per day (Figure 1, graph (a)). Fluctuations in traffic volume over short distances are influenced by various factors. Notably, there are considerably larger traffic volumes within Stjørdal to approximately 2.5 km towards Storlien. The European route E14 runs through the city center of Stjørdal, contributing to high traffic levels in the initial few kilometers.
The Norwegian Public Roads Administration (NPRA) recommends maintenance based on AADT values. High-traffic roads should be treated when rut depths exceed 20 mm, while low-traffic roads should be treated when rut depths exceed 25 mm. Adjusting AADT values based on road width is essential. Throughout the analyses, AADT values are expressed in ten-thousands.
2.2.4 Asphalt Type
Asphalt mixtures in cold-temperature regions like Norway are specifically designed to withstand harsh climatic conditions. These mixtures may include Polymer Modified Asphalt (PMA) to prevent cracking due to freeze-thaw cycles and low temperatures, and low-temperature performance grade binders to maintain flexibility at sub-zero temperatures. Dense-graded aggregates are used to minimize water infiltration, crucial for preventing freeze-thaw damage. The primary focus is on enhancing flexibility, low-temperature performance, and resistance to moisture, employing softer binders, dense-graded aggregates, and anti-stripping agents (Snilsberg et al. 2006, d’Angelo et al. 2008, Saba et al. 2009, Ødegård 2015).
Asphalt concrete (Ac), asphalt gravel concrete (Agc), and stone mastic asphalt (Sma) are the primary surface mixtures used in Norway. Sma, known for its superior wear resistance and deformation resistance due to the larger- and high-quality-stone in its composition, is typically employed on high-traffic roads. In contrast, Ac and Agc are commonly used on low-traffic roads, courtyards, sidewalks, and bicycle paths.
(a) AADT/Asphalt type and lane width
@@ 2.2.4 Asphalt Type
Figure 1: AADT together with asphalt type, and lane width (a), and rutting for a short stretch of E14 (for years 2019-2020, mm/year — missing in red) (b).
@@ 2.3 Exploratory analysis of rutting
On average, over all stretches and years, the overall rutting is $1.62$ mm/year; see Figure  2 (a). The density plots, Figure  2 graph (b), show rutting for all segments for four selected years (2017-2020). Some years having higher rutting levels and are more widely distributed than others. The lines in graphs (c) and (d) represent the empirical correlation of rutting values with respect to distance. Specifically, graph (c) shows the segment mean rutting over all years (see Section 2.2.1), and graph (d) displays the correlation for two selected years (2011-2012). These values are computed from the observed rutting at 20-meter segments, illustrating the empirical correlation as a function of the distance between the segments (Székely et al. 2007, Jentsch et al. 2020). Based on the data, we present results for distances up to 300 metres.
In graph (d), the x-axis represents the distance in meters, while the y-axis shows the calculated correlation between rutting values. This graph highlights the spatial dependencies of pavement rutting along the road for specific years. When compared to graph (c), which reveals long-term spatial patterns and trends, graph (d) demonstrates that the correlation of rutting values tends to diminish more quickly over shorter distances.
Graph (c) indicates that there are meaningful correlations in rutting values for distances up to approximately 150 meters, suggesting that these measurements capture underlying trends that persist over longer stretches of the road. In contrast, graph (d) shows that for specific years, the correlation between rutting values is important only up to about 100 meters. This shorter correlation distance implies that the rutting patterns observed in a given year are influenced by more localized factors and exhibit greater variability over shorter distances.
The difference in correlation distances between the two graphs shows the importance of considering both short-term and long-term spatial dependencies when analyzing pavement conditions. While long-term data provide a comprehensive view of how road conditions evolve, short-term data, as depicted in graph (d), highlight the influence of yearly variations and localized factors on pavement rutting. This dual approach enables a better understanding of the factors affecting road quality and can inform more effective maintenance and rehabilitation strategies.
(a) Observed rutting
s Refer to caption
(b) Density plot of rutting
(c) Autocorrelation of segment mean rutting
(d) Autocorrelation for years 2011-2012
Figure 2: Graph (a): Observed rutting for each year (2017-2020), the overall mean rutting ($\bar{r}=1.62$ mm/year (for all years, 2011-2020) (     ), and segment mean rutting ($\bar{r}_{i}$, over all 11 years) ( ). Graph (b): Density plots for rutting. Graphs (c) and (d) show how rutting values are correlated with distance. Spatial dependencies in rutting exist up to 150 metres.
@@ 3 Latent Gaussian models for Rutting
A Bayesian latent Gaussian model is a three-stage hierarchical model consisting of a likelihood, a latent Gaussian field, and hyperparameters. In this paper, the modelled rutting (i.e., all rutting behavior) at location $s_{i}$ in year $t$ is denoted by $\eta_{t}(s_{i})$ and it is modelled as a latent Gaussian field that includes explanatory variables and spatial effects. We specify six (6) models with different explanatory variables included and with and without spatial and random effects.
3.1 Likelihood: observed rutting given modelled rutting
We denote the observed rutting (detected and measured) for segment $i$ in year $t$ as $r_{i,t}$. The likelihood of the rutting, $r_{i,t}$, at road segment $i=1,...,n$ in year $t=1,...,T$ is expressed as